You are able to produce a home battery, that is inexpensive and easy to find, using tin cans as anode and copper wires as cathode and electrolyte of sodium chloride solution. In your poster, label the galvanic cell and computation using half-reactions from Standard Reduction Potentials at 25⁰C that this will create electricity.


1
Expert's answer
2021-01-11T04:44:54-0500

At Anode = Sn "\to" Sn+2 + 2e- ( E0 {Sn/Sn+2} = -0.14 )

At Cathode = Cu+2 + 2e- "\to" Cu ( E0 {Cu+2/Cu} = 0.34 )

Overall Net ionic equation =

Cu+2 + Sn "\to" Cu + Sn+2

E0 = 0.48

In the below diagram X = tin can and

Y = copper wires